OpenStudy (anonymous):

What is 80 ÷ (-50)? I got -1.6

OpenStudy (anonymous):

it has to be a fraction D:

OpenStudy (anonymous):

divide each number by 10

OpenStudy (anonymous):

you get \[\frac{8}{-5}=-\frac{8}{5}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

OK, so what is 8 divided by 5 with a remainder?

OpenStudy (anonymous):

has to be -1 then a fraction next to it

OpenStudy (anonymous):

SAME THING 1.6!!!

OpenStudy (anonymous):

i need a fraction!!!

OpenStudy (anonymous):

-8/5 is 1 with a remainder of -3, so \[-1\frac{3}{5}\]